# Verifiable Fine-Tuning for LLMs: Zero-Knowledge Training Proofs Bound to Data Provenance and Policy

### v2

@@ I Introduction
Watermarking techniques have also been advanced as a lightweight provenance signal for model outputs, but their reliability is contested. Classic and modern schemes embed statistical beacons into generated text and images [kirchenbauer2023watermark, wen2023treering], with recent work demonstrating scalable detectors for LLM text at industrial scale [dathathri2024natureWatermark]. Yet watermarks can be fragile or learnable by downstream models [gu2024learnability], susceptible to removal or theft [jovanovic2024wmstealing, zhang2024mipSteal], and often deliver weakest guarantees precisely where high-stakes provenance is needed, prompting more robust coding-theoretic designs [zhao2023provable, qu2025multibit] and broader surveys that map the design space and limitations [liu2024textwmSurvey]. These observations reinforce the case for training-time verifiability that does not rely solely on output beacons.
@@
In parallel, trusted execution and attestation are being used to produce verifiable “property cards” about ML artifacts, offering complementary hardware-backed assurances when cryptographic proofs are too costly or incomplete [duddu2025laminator]. However, TEEs alone cannot resolve disputes about which data were used or whether the optimizer trajectory conformed to policy, again pointing toward hybrid designs that combine attestations, provenance capture, and cryptographic proofs.
@@
The need for verifiable training is especially acute in federated and edge-centric learning, where heterogeneous clients, intermittent connectivity, and privacy constraints complicate governance. Practical systems already struggle with non-IID data, communication bottlenecks, and asynchrony [liu2021communication, liu2021adaptive, wang2021resource]. Algorithmic advances target adaptive regularization and knowledge distillation for robustness [liu2023adaptive], decentralized aggregation with layer-wise strategies [liu2023yoga], and personalization frameworks that better reflect client idiosyncrasies [yao2024ferrari, yan2024peaches]. Neural architecture search and progressive training mitigate heterogeneity in on-device settings [10251628, liu2024enhancing], while LoRA- and quantization-aware fine-tuning reduce resource footprints for practical deployments [li2025fedquad]. Recent work explores probabilistic communication and local-update strategies to accelerate decentralized training under diverse edges [liu2025accelerating, liu2025adaptive], pipelines for collaborative end–cloud inference [gao2025accelerating], strategies for MoE inference efficiency [yan2025accelerating], communication-efficient decentralized graph learning [wang2025towards], and mitigating catastrophic forgetting during federated fine-tuning [huo2025mitigating]. These directions improve scalability but simultaneously enlarge the attack surface and the number of untrusted actors—prime territory for verifiable fine-tuning protocols that certify data usage and update correctness [schlegel2025isProvenance, zkvml2023survey].
@@
Ubiquitous sensing further complicates the risk calculus. Wi-Fi and RF modalities, often intertwined with intelligent assistants, have been shown to support keystroke eavesdropping from commodity microphones [10587029], undetectable attacks on PHY-layer fingerprint-based authentication [10339891], and open-set gesture and attention-based recognition that blur lines between ambient intelligence and surveillance [10899398, 10090421]. Multimodal affective computing demonstrates how Wi-Fi CSI and vision together can unobtrusively infer emotion, expanding the privacy stakes of data pipelines [10149418]. On the defensive and robust-sensing side, prior work models and mitigates co-channel interference in activity recognition through subcarrier correlation selection and interference-independent phase components [9075376, 9613773]. Healthcare-oriented Wi-Fi sensing shows both potential and sensitivity, from target-oriented respiratory monitoring to pulmonary function estimation without physical contact [10804189, 10699348]. RF-based writing and identity inference illustrates how ordinary RFID can leak highly personal behaviors [feng2025rf]. Even adjacent work on label-noise suppression and distributionally robust facial expression recognition underscores how data curation and reliability shape downstream behaviors that later demand auditing [10916982, cui2025learningheterogeneitygeneralizingdynamic]. These examples motivate verifiable training pipelines that can convincingly demonstrate which signals were included, which were excluded, and how updates adhered to declared policies [schelter2024provenanceScreen].
@@
Our perspective in this paper is that verifiable fine-tuning should be treated as a first-class systems problem that composes three pillars: (i) cryptographic protocols that prove the existence of a legitimate gradient-based path from a public initialization to the released model under declared hyperparameters [sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm], (ii) provenance capture that binds those proofs to accountable, inspectable data and configuration claims [schlegel2023mlflow2prov, schlegel2025isProvenance, nikolic2025provenance], and (iii) audit mechanisms that recognize the limits of post-hoc inference and instead emphasize pre-commitment and incentive-compatible verification [jia2021proof, fang2022limits, zhao2024polIncentive]. We position our approach within this emerging landscape and focus on end-to-end verifiability for fine-tuning, complementing watermarking-based output attribution [kirchenbauer2023watermark, dathathri2024natureWatermark] and hardware-assisted attestations [duddu2025laminator], while remaining deployable in federated and edge scenarios [10251628, yao2024ferrari].
@@ II-C Output Attribution via Watermarking and Its Limits
Watermarks are attractive because they are lightweight and retrofittable. Early methods for LLM text embed pseudo-random beacons that can be detected statistically without changing human-perceived fluency [kirchenbauer2023watermark]. Large-scale evaluations suggest that scalable detectors can operate at Internet scale, though operating points depend sensitively on sampling and domain shifts [dathathri2024natureWatermark]. For images, invisible but robust fingerprinting for diffusion models extends the concept to the visual domain [wen2023treering]. Yet the reliability of such signals is contested: formal analyses show regimes where watermarks are learnable by downstream models or where adversaries can manipulate generation to evade detection [gu2024learnability], while empirical attacks demonstrate removal and even theft of watermark keys [jovanovic2024wmstealing, zhang2024mipSteal]. This tension has spurred coding-theoretic and cryptographic redesigns that seek provable robustness guarantees under realistic edits [zhao2023provable, qu2025multibit], as well as surveys that summarize the trade space and caveats for deployment [liu2024textwmSurvey]. In our scope, watermarking serves as an output-side signal that complements, but cannot replace, training-time verifiability.
@@ II-D Hardware-Assisted Attestation and ML Property Cards
Trusted execution environments (TEEs) and attestation mechanisms promise stronger integrity for training and deployment pipelines by anchoring code identity and configuration in hardware-rooted proofs. Recent work moves beyond attest-or-prove dichotomies and advocates verifiable property cards: structured attestations about what was trained and how, designed to be composable with cryptographic verification and external audits [duddu2025laminator]. In practice, TEEs can justify sensitive steps (e.g., decrypting private data, enforcing configuration constraints), but evidence of correct gradient updates and data usage still benefits from VML-style proof artifacts. We therefore treat TEEs as trustworthy enforcers around data access and hyperparameter policy, and ZK proofs as trustworthy witnesses of the numerical computations themselves.
@@ II-E Federated, Decentralized, and Edge Learning
Much of the urgency for verifiable fine-tuning arises in federated and decentralized training, where a large, heterogeneous population of clients interact under limited observability. Classical bottlenecks involve communication costs, asynchrony, and non-IID data; communication-efficient asynchronous protocols and hierarchical aggregation address some of these constraints and remain widely studied [liu2021communication, wang2021resource, liu2021adaptive]. Robustness and accuracy under heterogeneity have been pursued via adaptive regularization and knowledge distillation for federated updates [liu2023adaptive], decentralized layer-wise aggregation that aligns model structure with communication budgets [liu2023yoga], and personalization frameworks that better capture client idiosyncrasies and constraints [yao2024ferrari, yan2024peaches]. Neural architecture search at the federated edge and progressive training offer pathways to efficient on-device learning in non-IID regimes [10251628, liu2024enhancing].
Recent work also explores the dynamics of where and how to expend scarce resources. Experience-driven model migration considers heterogeneous edge networks with deep reinforcement learning to decide which models should move where, trading off accuracy and bandwidth [10508334]. For more demanding model classes, layer-wise LoRA deployment and activation quantization enable federated fine-tuning with reduced footprint, moving beyond naïve parameter-efficient fine-tuning with uniform recipes [li2025fedquad]. In decentralized regimes, probabilistic communication strategies and adaptive local updates can accelerate convergence and reduce staleness sensitivity [liu2025accelerating, liu2025adaptive]. Pushing to inference-time optimizations, pipeline-level decisions for end–cloud collaborative inference reduce bubbles and improve device utilization [gao2025accelerating], while MoE-specific strategies decrease routing overhead during on-device or near-edge inference [yan2025accelerating]. The federated graph setting adds its own twist: communication-efficient decentralized learning over non-IID graph-structured data requires new aggregation and sampling strategies [wang2025towards]. Catastrophic forgetting during federated fine-tuning is another stumbling block; adaptive block expansion in transformers ameliorates forgetting without ballooning communication [huo2025mitigating].
@@ II-F Security, Privacy, and Sensing Contexts that Raise the Stakes
Application contexts that blend ambient sensing with learning heighten the need for transparency about what signals enter training. Empirical studies show that commodity devices can support powerful side-channel inferences: smartphone microphones can eavesdrop on keystrokes with notable robustness across environments and typing habits [10587029]; attacks on PHY-layer fingerprint-based Wi-Fi authentication reveal that signal-level forgeries can evade detectors, challenging assumed trust anchors [10339891]. On the recognition side, Wi-Fi CSI has enabled attention-based gesture recognition and robust open-set gesture frameworks that work without special wearables, which in turn amplifies the sensitivity of datasets used for training [10090421, 10899398]. Multimodal affective computing compounds the effect: fusing Wi-Fi and vision allows unobtrusive inference of emotions via gestures and facial cues, with ramifications for consent and data minimization [10149418].
Defensive methodology exists. Anti-interference activity recognition can be improved by selecting subcarriers and modeling correlation structures, and interference-independent phase components reduce co-channel noise, both of which speak to the engineering burden of robust data curation before training [9075376, 9613773]. In healthcare, Wi-Fi-based respiratory monitoring and even contactless pulmonary function estimation showcase socially valuable use cases, but these also underline why verifiable declarations about data scope and preprocessing matter [10804189, 10699348]. Finally, commodity RFID-based systems have shown the feasibility of inferring handwriting content and user identity at a distance, again foregrounding how seemingly innocuous telemetry leaks highly personal attributes [feng2025rf]. For our purposes, these domains supply concrete scenarios where VFT can demonstrate, for example, that no raw audio from microphones was included, or that training included only synthetically generated CSI segments with declared distributions, binding those claims to cryptographic commitments.
@@ II-H Putting It Together: A Holistic View
The prior art suggests that no single technique suffices. Watermarks and model provenance tests address outputs and released artifacts [kirchenbauer2023watermark, liu2024textwmSurvey, nikolic2025provenance], but they struggle with adversarial transformations and lack formal guarantees in high-stakes settings [gu2024learnability, jovanovic2024wmstealing, zhang2024mipSteal]. Provenance capture and distributional attestation strengthen claims about input data [schelter2024provenanceScreen, schlegel2025isProvenance, dold2024attesting], yet without computation-level evidence they cannot exclude subtle deviations, such as skipped optimizer steps or modified learning-rate schedules. TEEs provide hardware roots of trust for code identity and configuration and can furnish property cards, but they offer limited transparency about the numerical content of large-scale training [duddu2025laminator]. Zero-knowledge and succinct proof systems, in turn, can certify computations but must be engineered to be affordable and compatible with ML numerics [sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm, taylorZKML2025].
In distributed settings, the gap widens: a federated organizer must balance communication, personalization, and fairness while defending against strategic participants [yao2024ferrari, yan2024peaches, liu2023adaptive]. Practical deployments need resource-aware methods—experience-driven migration, LoRA- and quantization-aware fine-tuning, decentralized aggregation, and asynchronous communication—to keep costs manageable [10508334, li2025fedquad, liu2025accelerating, liu2025adaptive, liu2023yoga, liu2021communication]. Modern pipelines should also acknowledge operational realities such as end–cloud collaboration, MoE routing, graph signals, and continual learning, aligning verification granularity with system bottlenecks [gao2025accelerating, yan2025accelerating, wang2025towards, huo2025mitigating, 10251628, liu2024enhancing]. The federated literature even starts to incorporate explicit verifiability in data contribution protocols and to explore probabilistic audit designs suited to decentralized LLMs [zhu2024risefl, probAudit2025].
Finally, application domains featuring ambient RF and Wi-Fi sensing vividly illustrate the stakes: side channels and ambient inference expand the privacy surface, while medical and affective applications raise the bar for compliance and governance [10587029, 10339891, 10090421, 10899398, 10149418, 9075376, 9613773, 10804189, 10699348, feng2025rf]. In such contexts, an end-to-end verifiable fine-tuning pipeline can bridge the trust gap by committing to the scope and transformations of sensitive signals and by providing cryptographic evidence that numerical updates followed declared policies. This paper builds on these insights and contributes a protocol and system architecture for verifiable fine-tuning that composes cryptographic proofs, provenance capture, and audit mechanisms into a deployable workflow.
@@ II-I Scope and Differentiation
Within this landscape, our work differs from prior ZK-ML for training in three ways. First, we explicitly bind proof objects to provenance claims, rather than treating data selection as orthogonal pre-processing. This design enables verifiers to audit both which batches were admissible and how they were used, marrying the strengths of distributional attestation and lineage capture with computation proofs [dold2024attesting, schelter2024provenanceScreen, schlegel2025isProvenance]. Second, we focus on parameter-efficient fine-tuning regimes common in real deployments—LoRA and post-training adapters—so that proof sizes scale with adapted subspaces, while still covering optimizer semantics and numerical approximations consistent with recent ZK-LLM advances [sun2024zkllm, taylorZKML2025]. Third, we design partial, round-based verification hooks compatible with federated and decentralized orchestration, benefiting from insights in communication-efficient and personalized learning [yao2024ferrari, yan2024peaches, liu2025accelerating]. Our evaluation examines both verification cost and end-task quality, acknowledging practical constraints emphasized by systems studies on pipeline scheduling and edge collaboration [gao2025accelerating, 10251628, liu2024enhancing]. We regard watermarking and model provenance testing as complementary post-release signals and discuss how our commitments can assist downstream detectors by making output attribution less ambiguous [kirchenbauer2023watermark, nikolic2025provenance, liu2024textwmSurvey].

### v3

@@ I Introduction
Watermarking techniques have also been advanced as a lightweight provenance signal for model outputs, but their reliability is contested. Classic and modern schemes embed statistical beacons into generated text and images [kirchenbauer2023watermark, wen2023treering], with recent work demonstrating scalable detectors for LLM text at industrial scale [dathathri2024natureWatermark]. Yet watermarks can be fragile or learnable by downstream models [gu2024learnability], susceptible to removal or theft [jovanovic2024wmstealing, zhang2024mipSteal], and often deliver weakest guarantees precisely where high-stakes provenance is needed, prompting more robust coding-theoretic designs [qu2025multibit] and broader surveys that map the design space and limitations [liu2024textwmSurvey]. These observations reinforce the case for training-time verifiability that does not rely solely on output beacons.
@@
In parallel, trusted execution and attestation are being used to produce verifiable “property cards” about ML artifacts, offering complementary hardware-backed assurances when cryptographic proofs are too costly or incomplete. However, TEEs alone cannot resolve disputes about which data were used or whether the optimizer trajectory conformed to policy, again pointing toward hybrid designs that combine attestations, provenance capture, and cryptographic proofs.
@@
The need for verifiable training is especially acute in federated and edge-centric learning, where heterogeneous clients, intermittent connectivity, and privacy constraints complicate governance. Practical systems already struggle with non-IID data, communication bottlenecks, and asynchrony. Algorithmic advances target adaptive regularization and knowledge distillation for robustness, decentralized aggregation with layer-wise strategies, and personalization frameworks that better reflect client idiosyncrasies . Neural architecture search and progressive training mitigate heterogeneity in on-device settings , while LoRA- and quantization-aware fine-tuning reduce resource footprints for practical deployments. Recent work explores probabilistic communication and local-update strategies to accelerate decentralized training under diverse edges , pipelines for collaborative end–cloud inference, strategies for MoE inference efficiency, communication-efficient decentralized graph learning, and mitigating catastrophic forgetting during federated fine-tuning. These directions improve scalability but simultaneously enlarge the attack surface and the number of untrusted actors—prime territory for verifiable fine-tuning protocols that certify data usage and update correctness [schlegel2025isProvenance, zkvml2023survey].
@@
Ubiquitous sensing further complicates the risk calculus. Wi-Fi and RF modalities, often intertwined with intelligent assistants, have been shown to support keystroke eavesdropping from commodity microphones, undetectable attacks on PHY-layer fingerprint-based authentication, and open-set gesture and attention-based recognition that blur lines between ambient intelligence and surveillance . Multimodal affective computing demonstrates how Wi-Fi CSI and vision together can unobtrusively infer emotion, expanding the privacy stakes of data pipelines. On the defensive and robust-sensing side, prior work models and mitigates co-channel interference in activity recognition through subcarrier correlation selection and interference-independent phase components . Healthcare-oriented Wi-Fi sensing shows both potential and sensitivity, from target-oriented respiratory monitoring to pulmonary function estimation without physical contact . RF-based writing and identity inference illustrates how ordinary RFID can leak highly personal behaviors. Even adjacent work on label-noise suppression and distributionally robust facial expression recognition underscores how data curation and reliability shape downstream behaviors that later demand auditing . These examples motivate verifiable training pipelines that can convincingly demonstrate which signals were included, which were excluded, and how updates adhered to declared policies [schelter2024provenanceScreen].
@@
Our perspective in this paper is that verifiable fine-tuning should be treated as a first-class systems problem that composes three pillars: (i) cryptographic protocols that prove the existence of a legitimate gradient-based path from a public initialization to the released model under declared hyperparameters [sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm], (ii) provenance capture that binds those proofs to accountable, inspectable data and configuration claims [schlegel2023mlflow2prov, schlegel2025isProvenance, nikolic2025provenance], and (iii) audit mechanisms that recognize the limits of post-hoc inference and instead emphasize pre-commitment and incentive-compatible verification [jia2021proof, fang2022limits, zhao2024polIncentive]. We position our approach within this emerging landscape and focus on end-to-end verifiability for fine-tuning, complementing watermarking-based output attribution [kirchenbauer2023watermark, dathathri2024natureWatermark] and hardware-assisted attestations, while remaining deployable in federated and edge scenarios .
@@ II-C Output Attribution via Watermarking and Its Limits
Watermarks are attractive because they are lightweight and retrofittable. Early methods for LLM text embed pseudo-random beacons that can be detected statistically without changing human-perceived fluency [kirchenbauer2023watermark]. Large-scale evaluations suggest that scalable detectors can operate at Internet scale, though operating points depend sensitively on sampling and domain shifts [dathathri2024natureWatermark]. For images, invisible but robust fingerprinting for diffusion models extends the concept to the visual domain [wen2023treering]. Yet the reliability of such signals is contested: formal analyses show regimes where watermarks are learnable by downstream models or where adversaries can manipulate generation to evade detection [gu2024learnability], while empirical attacks demonstrate removal and even theft of watermark keys [jovanovic2024wmstealing, zhang2024mipSteal]. This tension has spurred coding-theoretic and cryptographic redesigns that seek provable robustness guarantees under realistic edits [qu2025multibit], as well as surveys that summarize the trade space and caveats for deployment [liu2024textwmSurvey]. In our scope, watermarking serves as an signal that complements, but cannot replace, training-time verifiability.
@@ II-D Hardware-Assisted Attestation and ML Property Cards
Trusted execution environments (TEEs) and attestation mechanisms promise stronger integrity for training and deployment pipelines by anchoring code identity and configuration in hardware-rooted proofs. Recent work moves beyond attest-or-prove dichotomies and advocates verifiable property cards: structured attestations about what was trained and how, designed to be composable with cryptographic verification and external audits. In practice, TEEs can justify sensitive steps (e.g., decrypting private data, enforcing configuration constraints), but evidence of correct gradient updates and data usage still benefits from VML-style proof artifacts. We therefore treat TEEs as trustworthy enforcers around data access and hyperparameter policy, and ZK proofs as trustworthy witnesses of the numerical computations themselves.
@@ II-E Federated, Decentralized, and Edge Learning
Much of the urgency for verifiable fine-tuning arises in federated and decentralized training, where a large, heterogeneous population of clients interact under limited observability. Classical bottlenecks involve communication costs, asynchrony, and non-IID data; communication-efficient asynchronous protocols and hierarchical aggregation address some of these constraints and remain widely studied . Robustness and accuracy under heterogeneity have been pursued via adaptive regularization and knowledge distillation for federated updates, decentralized layer-wise aggregation that aligns model structure with communication budgets, and personalization frameworks that better capture client idiosyncrasies and constraints . Neural architecture search at the federated edge and progressive training offer pathways to efficient on-device learning in non-IID regimes .
Recent work also explores the dynamics of where and how to expend scarce resources. Experience-driven model migration considers heterogeneous edge networks with deep reinforcement learning to decide which models should move where, trading off accuracy and bandwidth. For more demanding model classes, layer-wise LoRA deployment and activation quantization enable federated fine-tuning with reduced footprint, moving beyond naïve parameter-efficient fine-tuning with uniform recipes. In decentralized regimes, probabilistic communication strategies and adaptive local updates can accelerate convergence and reduce staleness sensitivity . Pushing to inference-time optimizations, pipeline-level decisions for end–cloud collaborative inference reduce bubbles and improve device utilization, while MoE-specific strategies decrease routing overhead during on-device or near-edge inference. The federated graph setting adds its own twist: communication-efficient decentralized learning over non-IID graph-structured data requires new aggregation and sampling strategies. Catastrophic forgetting during federated fine-tuning is another stumbling block; adaptive block expansion in transformers ameliorates forgetting without ballooning communication.
@@ II-F Security, Privacy, and Sensing Contexts that Raise the Stakes
Application contexts that blend ambient sensing with learning heighten the need for transparency about what signals enter training. Empirical studies show that commodity devices can support powerful side-channel inferences: smartphone microphones can eavesdrop on keystrokes with notable robustness across environments and typing habits; attacks on PHY-layer fingerprint-based Wi-Fi authentication reveal that signal-level forgeries can evade detectors, challenging assumed trust anchors. On the recognition side, Wi-Fi CSI has enabled attention-based gesture recognition and robust open-set gesture frameworks that work without special wearables, which in turn amplifies the sensitivity of datasets used for training . Multimodal affective computing compounds the effect: fusing Wi-Fi and vision allows unobtrusive inference of emotions via gestures and facial cues, with ramifications for consent and data minimization.
Defensive methodology exists. Anti-interference activity recognition can be improved by selecting subcarriers and modeling correlation structures, and interference-independent phase components reduce co-channel noise, both of which speak to the engineering burden of robust data curation before training . In healthcare, Wi-Fi-based respiratory monitoring and even contactless pulmonary function estimation showcase socially valuable use cases, but these also underline why verifiable declarations about data scope and preprocessing matter . Finally, commodity RFID-based systems have shown the feasibility of inferring handwriting content and user identity at a distance, again foregrounding how seemingly innocuous telemetry leaks highly personal attributes. For our purposes, these domains supply concrete scenarios where VFT can demonstrate, for example, that no raw audio from microphones was included, or that training included only synthetically generated CSI segments with declared distributions, binding those claims to cryptographic commitments.
@@ II-H Putting It Together: A Holistic View
The prior art suggests that no single technique suffices. Watermarks and model provenance tests address outputs and released artifacts [kirchenbauer2023watermark, liu2024textwmSurvey, nikolic2025provenance], but they struggle with adversarial transformations and lack formal guarantees in high-stakes settings [gu2024learnability, jovanovic2024wmstealing, zhang2024mipSteal]. Provenance capture and distributional attestation strengthen claims about input data [schelter2024provenanceScreen, schlegel2025isProvenance, dold2024attesting], yet without computation-level evidence they cannot exclude subtle deviations, such as skipped optimizer steps or modified learning-rate schedules. TEEs provide hardware roots of trust for code identity and configuration and can furnish property cards, but they offer limited transparency about the numerical content of large-scale training. Zero-knowledge and succinct proof systems, in turn, can certify computations but must be engineered to be affordable and compatible with ML numerics [sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm, taylorZKML2025].
In distributed settings, the gap widens: a federated organizer must balance communication, personalization, and fairness while defending against strategic participants . Practical deployments need resource-aware methods—experience-driven migration, LoRA- and quantization-aware fine-tuning, decentralized aggregation, and asynchronous communication—to keep costs manageable . Modern pipelines should also acknowledge operational realities such as end–cloud collaboration, MoE routing, graph signals, and continual learning, aligning verification granularity with system bottlenecks . The federated literature even starts to incorporate explicit verifiability in data contribution protocols and to explore probabilistic audit designs suited to decentralized LLMs [zhu2024risefl, probAudit2025].
Finally, application domains featuring ambient RF and Wi-Fi sensing vividly illustrate the stakes: side channels and ambient inference expand the privacy surface, while medical and affective applications raise the bar for compliance and governance . In such contexts, an end-to-end verifiable fine-tuning pipeline can bridge the trust gap by committing to the scope and transformations of sensitive signals and by providing cryptographic evidence that numerical updates followed declared policies. This paper builds on these insights and contributes a protocol and system architecture for verifiable fine-tuning that composes cryptographic proofs, provenance capture, and audit mechanisms into a deployable workflow.
@@ II-I Scope and Differentiation
Within this landscape, our work differs from prior ZK-ML for training in three ways. First, we explicitly bind proof objects to provenance claims, rather than treating data selection as orthogonal pre-processing. This design enables verifiers to audit both which batches were admissible and how they were used, marrying the strengths of distributional attestation and lineage capture with computation proofs [dold2024attesting, schelter2024provenanceScreen, schlegel2025isProvenance]. Second, we focus on parameter-efficient fine-tuning regimes common in real deployments—LoRA and post-training adapters—so that proof sizes scale with adapted subspaces, while still covering optimizer semantics and numerical approximations consistent with recent ZK-LLM advances [sun2024zkllm, taylorZKML2025]. Third, we design partial, round-based verification hooks compatible with federated and decentralized orchestration, benefiting from insights in communication-efficient and personalized learning . Our evaluation examines both verification cost and end-task quality, acknowledging practical constraints emphasized by systems studies on pipeline scheduling and edge collaboration . We regard watermarking and model provenance testing as complementary post-release signals and discuss how our commitments can assist downstream detectors by making output attribution less ambiguous [kirchenbauer2023watermark, nikolic2025provenance, liu2024textwmSurvey].